OpenStudy (anonymous):

Factor out the Monomial 1/(2x^2+4x)= 1/x(

OpenStudy (jhannybean):

For this one you have \[\large \frac{1}{2x^2+4x}\] There is a common multiplier at the bottom,can you tell me what it is?

OpenStudy (jhannybean):

hint: it's2 terms,and one of them is a multiple of 2.

OpenStudy (anonymous):

Well I see 1/2x being taken out.

OpenStudy (jhannybean):

Good! not "taken out" per say,but "factored out" \[\large \frac{1}{2x(x+2)}\]

OpenStudy (anonymous):

The problem only asks to take out (1/x). I'm thinking that I write it like 1/x(1/(2x+4) Does that look correct?

OpenStudy (jhannybean):

Well,if you took out only 1 x from the denominator...you would have.. \[\large \frac{1}{x(2x+2)}\]

OpenStudy (jhannybean):

ack,sorry. 4 instead of 2.

OpenStudy (anonymous):

Alright I think that answers my question thanks :)

OpenStudy (jhannybean):

\[\large \frac{1}{x(2x+4)}\]

OpenStudy (jhannybean):

Which is kind of incorrect because when you're "taking out" or"factoring" a function, you look for common multipliers that appear in all terms, and take out the LCM. that includes numbers and variables.

OpenStudy (jhannybean):

Between 2 and 4, 2 is the LCM, and between x and x^(2), x is the LCM.